# VidLaDA: Bidirectional Diffusion Large Language Models for Efficient Video Understanding

### v1

@@ Abstract
Standard Autoregressive Video LLMs inevitably suffer from causal masking biases that hinder global spatiotemporal modeling, leading to suboptimal understanding efficiency. We propose VidLaDA, a Video LLM based on Diffusion Language Model utilizing bidirectional attention to capture bidirectional dependencies. To further tackle the inference bottleneck of diffusion decoding on massive video tokens, we introduce MARS-Cache. This framework accelerates inference by combining asynchronous visual cache refreshing with frame-wise chunk attention, effectively pruning redundancy while preserving global connectivity via anchor tokens. Extensive experiments show VidLaDA outperforms diffusion baselines and rivals state-of-the-art autoregressive models (e.g., Qwen2.5-VL and LLaVA-Video), with MARS-Cache delivering over 12× speedup without compromising reasoning accuracy. Code and checkpoints are open-sourced at https://github.com/ziHoHe/VidLaDA.
@@
Figure 1: The overall architecture of VidLaDA. Input video frames are encoded and spatially pooled (via $2\times 2$ downsampling) before being unrolled into a sequence of Spatiotemporal Visual Tokens $V$. These tokens, combined with the text prompt $P$ and the noised answer $X_{t}$, are processed by the Diffusion Language Model. Unlike autoregressive models, VidLaDA utilizes full bidirectional attention. This design enables global, unconstrained interactions both within and across visual and textual modalities, ultimately facilitating the parallel prediction of the target answer $X_{0}$.
@@ 1 Introduction
Video Large Language Models (Video LLMs) (Lin et al., 2024; Cheng et al., 2024) have recently emerged as a key paradigm for connecting visual perception with language reasoning, showing strong potential in tasks such as video captioning, spatiotemporal question answering, and long-horizon decision making. Most existing systems follow a standard recipe that couples a pretrained vision encoder (e.g., ViT (Zhu et al., 2023; Tschannen et al., 2025)) with an Autoregressive (AR) LLM (Grattafiori et al., 2024; Team et al., 2024). Video frames are first encoded into visual tokens and projected into the language embedding space, after which the model generates responses via next-token prediction under a causal attention mask (Lin et al., 2024).
@@
Moreover, video understanding is increasingly shifting from short-horizon perception to complex spatiotemporal reasoning (Ye et al., 2025b; Liu et al., 2025; Feng et al., 2025). Compared to images, videos are both temporally redundant and dynamically evolving, which often requires models to efficiently aggregate and understand key spatiotemporal evidence across the video and form a long-form reasoning chain. However, the standard AR paradigm faces a dual efficiency bottleneck in this context. First, regarding understanding efficiency (modeling front), strictly unidirectional temporal modeling can hinder capturing global temporal relationships across different moments in a video (Guo et al., 2025b; Chen et al., 2024b), which limits the effective aggregation of spatiotemporal evidence. Second, regarding generation efficiency (inference), AR decoding is inherently serial, making latency scale linearly with generated tokens and limiting generation efficiency for long-form reasoning for complex spatiotemporal reasoning. Conversely, Diffusion Language Models (DLMs) offer a compelling alternative: generation is formulated as iterative denoising in a discrete space with full bidirectional attention (Nie et al., 2025), which removes the causal constraint and enables parallel prediction of multiple tokens per step (Wu et al., 2025). This bidirectional decoding is advantageous for video understanding, where key spatiotemporal evidence should aggregate visual context globally across all frames and patches rather than follow a forced causal path.
@@
Figure 2: Comparison of Spatiotemporal Robustness. (a) Performance vs. spatial location of high-norm tokens. DLM-Based VLM remains invariant, whereas AR degrades when salient features shift from the start. (b) Performance vs. temporal location of the key event. DLM-Based VLM demonstrates stability across the timeline, while AR baselines show significant volatility. (c) DLM-Based VLM maintains high accuracy with fewer frames, demonstrating superior aggregation of sparse evidence compared to AR models.
Despite these advantages, directly applying DLMs to long-form video introduces a new efficiency bottleneck. Unlike AR inference, where prefix key/value states can be cached, and only the newest token is processed, vanilla DLM inference recomputes bidirectional attention over the entire multimodal sequence at every decoding step. With thousands of video tokens, this results in prohibitive $\mathcal{O}(|V|^{2})$ attention cost per step, which can negate the speed benefits of parallel decoding. Moreover, generic DLM acceleration techniques developed for text-only settings (Wu et al., 2025) ignore multimodal heterogeneity: in our analysis, visual and textual states exhibit markedly different decoding dynamics, and inter-frame attention presents strong local temporal structure with a small set of globally influential anchor tokens. Treating all modalities and all layers identically, therefore wastes computation on repeatedly recomputing stable context, especially for video inputs.
To unlock bidirectional diffusion for efficient video understanding, we propose VidLaDA (Video-Language Diffusion with mAsking), a Video LLM built on the DLM backbone with bidirectional attention to maximize understanding efficiency, and MARS-Cache (Multi-modal Asynchronous Refreshing Strategy), an inference framework that prunes redundancy specific to multimodal diffusion to maximize generation efficiency. Trained via a multi-stage curriculum on a comprehensive video dataset incorporating our newly constructed collection, VidLaDA couples a strong vision encoder with DLM-based bidirectional decoding, enabling unconstrained global interactions across video tokens, prompts, and partially denoised answers. MARS-Cache accelerates long-video DLM inference by combining (i) frame-wise chunk attention to exploit temporal locality, reducing the visual refresh cost to linear $\mathcal{O}(|V|)$, (ii) adaptive anchor token searching to preserve necessary global connectivity, and (iii) modality- and depth-wise asynchronous cache refreshing to update stable visual context less frequently than dynamic textual states. Our main contributions are summarized as follows:
@@
    We propose MARS-Cache, a Multi-modal Asynchronous Refreshing Strategy, motivated by observations of modality-wise stability and attention structure during decoding. Substantially, MARS-Cache achieves a more than $12\times$ throughput improvement compared to vanilla DLM without compromising reasoning accuracy.
@@ 2 Preliminary
Problem Formulation. Given a video input $\mathcal{V}$ and a textual prompt $P$, the goal of a Video Large Language Model is to generate a target response $R=\{r_{1},r_{2},\dots,r_{L}\}$. The video is typically processed into visual tokens $V$, which form the multimodal context alongside prompt textual tokens $P$.
@@
In contrast, our VidLaDA is built upon Diffusion Language Models (DLMs) (Nie et al., 2025; Ye et al., 2025c), which treat generation as a bidirectional iterative denoising process. Instead of sequential prediction, DLMs utilize a Masked Diffusion Model (MDM) framework. During training, a subset of response tokens is randomly replaced by a special [MASK] token based on a timestep $t$. The model learns to predict the original identities of these masked tokens simultaneously (Nie et al., 2025). Moreover, the DLM architecture allows VidLaDA to employ a full bidirectional attention mechanism without causal masking. Consequently, the aggregation and prediction of any token depends on the entire global context (both unmasked and masked tokens in $V$, $P$, and $R$), facilitating unconstrained interaction between visual and textual modalities. For detailed mathematical formulations of the forward and reverse diffusion processes, please refer to Appendix A.
@@
Figure 3: Visualization of Attention Patterns. We display the attention score matrices across different denoising steps and layers. The heatmaps reveal two distinct structural properties utilized by our MARS-Cache design: (1) Chunk-wise Locality, visible as diagonal blocks where tokens primarily attend to their temporal neighbors, and (2) Global Anchor Tokens, manifested as prominent vertical bands where specific tokens consistently attract global attention from the entire sequence, regardless of the diffusion step or network depth.
@@ 3.1 Why Bidirectional DLM for Video Understanding?
The AR paradigm inherently limits the utilization of omnidirectional spatiotemporal features encoded by ViT, resulting in a strictly lower mutual information upper bound compared to a bidirectional decoding paradigm.
$\sum_{t=1}^{T}\sup I(z_{t}^{\text{bidi}};\mathcal{V})>\sum_{t=1}^{T}\sup I(z_{t}^{\text{uni}};\mathcal{V}),$ (1)
where $\mathcal{V}$ denotes the original visual inputs, and $z^{*}_{t}$ denotes the representation encoded by ViT and LLM at timestep $t$.
@@ 3.1.1 The Impact of Token Spatial Ordering
In ViT, the patch order is conventionally rasterized (top-left to bottom-right). Inspired by (Luo et al., 2025), we use the $\ell_{2}$-norm of the ViT output features as a proxy for the information density of visual tokens. We identify the top-$k$ high-norm tokens and virtually relocate them within the sequence fed to the LLM, but maintaining the same positional IDs in sequence. This design isolates the effect of causal visibility in the decoder from changes in spatial positional encoding. We define a position ratio $r\in[0,1]$, where $r=0$ places high-norm tokens at the start of the visual sequence and $r=1$ pushes them to the end.
As illustrated in Figure 2(a), we evaluate the models on the MME (Yin et al., 2024) benchmark under this shuffling protocol (cf. Appendix F.2). The Bidirectional DLM-Based VLM (You et al., 2025) exhibits a near-constant performance profile (red line), demonstrating invariance to the causal position of semantic features.
Conversely, AR-based models (LLaVA-OneVision (Li et al., 2024a), LLaVA-1.5 (Liu et al., 2024b)) display performance degradation, particularly when high-information tokens are shifted towards the middle or end of the sequence. This corroborates Proposition 3.1: AR models over-rely on early tokens as a receptive field to maintain semantic interpretation. When salient features are displaced from the high-visibility start positions to later positions (where visibility is lower), the autoregressive attention mechanism fails to allocate sufficient attention mass to them, treating them as contextually less significant solely due to their temporal index.
@@ 3.1.2 The Impact of Event Position and Sparsity
To verify spatiotemporal robustness in realistic settings beyond artificial single-image shuffling, we evaluate on ReXTime (Chen et al., 2024a) (cf. Appendix F.3), where each question depends on a specific event segment. We bucket samples by the temporal position ratio of the ground-truth event and report accuracy under uniform frame sampling.
@@
Notably, LLaVA-Video (Zhang et al., 2024b) exhibits the most severe instability despite extensive video training compared to LLaVA-OneVision (Li et al., 2024a), suggesting that scaling data alone cannot remove the structural limitations imposed by the early receptive field of AR. This limitation becomes more evident when event evidence is sparse. As shown in Figure 2(c), AR models degrade more sharply as fewer event frames are available, indicating that answer-relevant information of the event is more likely to be dispersed during deep semantic abstraction once it is not continuously reinforced under causal attention. In contrast, the bidirectional DLM-based Video LLM maintains a near-flat accuracy profile across event positions (Figure 2(b)) and remains accurate even under sparse event frames (Figure 2(c)). This behavior is consistent with Proposition 3.2: bidirectional decoding allows answer-relevant representations to directly aggregate global spatiotemporal context across all frames. This mechanism enhances understanding efficiency and prevents the irreversible loss of mid-sequence event details during deep semantic abstraction, even under sparse evidence conditions. Furthermore, it inherently eliminates the visibility asymmetry that induces early receptive field in AR decoding (Proposition 3.1).
@@ 3.2 Training Pipeline
Data Composition. To address the scarcity of training data for video understanding beyond 3 minutes, we first curate a dataset emphasizing 2-30 minute videos. The pipeline involves (1) temporal stratification, (2) instruction synthesis via LLM, (3) text-only bias filtering, and (4) VLM consistency voting (See Appendix D.1.1 for details).
Multi-Stage Training. We also employ a multi-stage training strategy for VidLaDA (See Appendix D.1.2 for details), including (1) short-clip temporal pre-training, (2) temporal scaling warm-up, and (3) long-form video expansion.
@@ 3.3 Model Architecture
Bidirectional Diffusion Language Model. The core unit of VidLaDA is the Diffusion Language Model with the full bidirectional attention mechanism. This architecture addresses the limitations of traditional Autoregressive (AR) VLMs to some extent. Standard AR models rely on causal attention, which imposes a strict left-to-right dependency. This structure introduces the asymmetric receptive field, where early visual tokens are structurally prevented from attending to subsequent visual or textual context from user input, disrupting the global topology of the vision (See Section 3.1). In contrast, VidLaDA leverages a full bidirectional attention mechanism. This design eliminates the causal constraint, allowing every token, whether visual or textual, to attend to the holistic sequence simultaneously. This ensures that the vision encoder and language model are deeply coupled, preserving the integrity of dynamic video information.
@@ 4.1 Empirical Observations
We analyze the internal state evolution during the decoding process and identify four distinct observations: (1) Chunk-wise Locality with Global Anchor Tokens. As shown in Figure 3, inter-frame interactions primarily exhibit local dependencies. However, a specific subset of anchor tokens consistently attracts high attention scores globally. Crucially, we observe that the spatial positions of these anchor tokens remain stable across diffusion steps, indicating that the global information hubs are determined early in the inference process. Furthermore, we observe a hierarchical inclusion property: the set of anchor tokens in deeper layers is a subset of those in shallow layers ($\mathcal{I}_{\text{deep}}\subset\mathcal{I}_{\text{shallow}}$), adhering to a dense-to-sparse pyramid structure. (2) Modality-Dependent Drift. As shown in Figure 5, there is a notable disparity in hidden state evolution between modalities. Text hidden states exhibit higher temporal drift across steps compared to vision tokens. (3) Depth-Dependent Stability Variance. As shown in Figure 5, hidden state drift is not uniform across network depth. Shallow layers exhibit high stability, while deep layers show significant volatility. (4) Progressive Attention Sparsity. As shown in Figure 6, the attention distribution evolves through the network depth, transitioning from a near-uniform distribution in shallow layers to a highly peaked, sparse distribution in deep layers.
@@
Table 1: Comparison with SOTA Video LLMs across comprehensive benchmarks. The models are categorized into AR-Based and DLM-Based models. #S: Model Size (Parameters). #F: Number of input frames. Models marked with * denote the reproduced results.

### v2

@@ Abstract
[…]
@@
Figure 1: The overall architecture of VidLaDA. Input video frames $\mathcal{V}$} are encoded and spatially pooled (via $2\times 2$ downsampling) before being unrolled into a sequence of Spatiotemporal Visual Tokens ${\bm{E}}^{\mathcal{V}}$}. These tokens, combined with the text prompt $P$ and the noised response $R^{t}$, are processed by the Diffusion Language Model. Unlike autoregressive models, VidLaDA utilizes full bidirectional attention. This design enables global, unconstrained interactions both within and across visual and textual modalities, ultimately facilitating the parallel prediction of the target response $R^{0}$.
@@ 1 Introduction
Video Large Language Models (Video LLMs) (Lin et al., 2024; Cheng et al., 2024) have recently emerged as a key paradigm for connecting visual perception with language reasoning, showing strong potential in tasks like video captioning, spatiotemporal question answering, and long-horizon decision making. Most existing systems follow a standard recipe that couples a pretrained vision encoder (e.g., ViT (Zhu et al., 2023; Tschannen et al., 2025)) with an Autoregressive (AR) LLM (Grattafiori et al., 2024; Team et al., 2024). Video frames are first encoded into visual embeddings and projected into the language embedding space, after which the model generates responses via next-token prediction under a causal attention mask (Lin et al., 2024).
@@
Moreover, video understanding is increasingly shifting from short-horizon perception to complex spatiotemporal reasoning (Ye et al., 2025b; Liu et al., 2025; Feng et al., 2025). Compared to images, videos are both temporally redundant and dynamically evolving, which often requires models to efficiently aggregate and understand key spatiotemporal evidence across the video and form a long-form reasoning chain. However, the standard AR paradigm faces a dual efficiency bottleneck in this context. First, regarding understanding efficiency as analyzed above, strictly unidirectional temporal modeling can hinder capturing global temporal relationships across different moments in a video (Guo et al., 2025b; Chen et al., 2024b), which limits the effective aggregation of spatiotemporal evidence. Second, regarding generation efficiency, AR decoding is inherently serial, making latency scale linearly with generated tokens and limiting generation efficiency for long-form reasoning for complex spatiotemporal reasoning. Conversely, Diffusion Language Models (DLMs) offer a compelling alternative: generation is formulated as iterative denoising in a discrete space with full bidirectional attention (Nie et al., 2025), which removes the causal constraint and enables parallel prediction of multiple tokens per step (Wu et al., 2025). This bidirectional decoding is advantageous for video understanding, where key spatiotemporal evidence should aggregate visual context globally across all frames and patches rather than follow a forced causal path.
@@
Figure 2: Comparison of Spatiotemporal Robustness. (a) Performance vs. spatial location of high-norm tokens. DLM baseline remains invariant, whereas AR degrades when salient features shift from the start. (b) Performance vs. temporal location of the key event. DLM baseline demonstrates stability across the timeline, while AR baselines show significant volatility. (c) DLM baseline maintains high accuracy with fewer frames, demonstrating superior aggregation of sparse evidence compared to AR models.
Despite these advantages, directly applying DLMs to long-form video introduces a new efficiency bottleneck. Unlike AR inference, where prefix key/value states can be cached, and only the newest token is processed, vanilla DLM inference recomputes bidirectional attention over the entire multimodal sequence at every decoding step. With thousands of video tokens ${\bm{E}}^{\mathcal{V}}$}, this results in prohibitive $\mathcal{O}(|{{\bm{E}}^{\mathcal{V}}}|^{2})$ attention cost per step, which can negate the speed benefits of parallel decoding. Moreover, generic DLM acceleration techniques developed for text-only settings (Wu et al., 2025) ignore multimodal heterogeneity: in our analysis, visual and textual states exhibit markedly different decoding dynamics, and inter-frame attention presents strong local temporal structure with a small set of globally influential anchor tokens. Treating all modalities and all layers identically, therefore wastes computation on repeatedly recomputing stable context, especially for video inputs.
To unlock bidirectional diffusion for efficient video understanding, we propose VidLaDA (Video-Language Diffusion with mAsking), a Video LLM built on the DLM backbone with bidirectional attention to maximize understanding efficiency, and MARS-Cache (Multi-modal Asynchronous Refreshing Strategy), an inference framework that prunes redundancy specific to multimodal diffusion to maximize generation efficiency. Trained via a multi-stage curriculum on a comprehensive video dataset incorporating our newly constructed collection, VidLaDA couples a strong vision encoder with DLM-based bidirectional decoding, enabling unconstrained global interactions across video tokens, prompts, and partially denoised responses. MARS-Cache accelerates DLM-based Video LLM decoding by combining (i) frame-wise chunk attention to exploit temporal locality, reducing the visual refresh cost to linear $\mathcal{O}(|{{\bm{E}}^{\mathcal{V}}}|)$, (ii) adaptive anchor token searching to preserve necessary global connectivity, and (iii) modality- and depth-wise asynchronous cache refreshing to update stable visual context less frequently than dynamic textual states. Our main contributions are summarized as follows:
@@
[…]
@@ 2 Preliminary
Problem Formulation. Given a video input $\mathcal{V}$} and a textual prompt $P$, the goal of a Video Large Language Model is to generate a target response $R=\{R_{1},\dots,R_{|R|}\}$. The video is typically processed into visual tokens ${\bm{E}}^{\mathcal{V}}$}, which form the multimodal context alongside prompt textual tokens $P$.
@@
In contrast, our VidLaDA is built upon Diffusion Language Models (DLMs) (Nie et al., 2025; Ye et al., 2025c), which treat generation as a bidirectional iterative denoising process. Instead of sequential prediction, DLMs utilize a Masked Diffusion Model (MDM) framework. During training, a subset of response tokens is randomly replaced by a special [MASK] token based on a timestep $t$. The model learns to predict the original identities of these masked tokens simultaneously (Nie et al., 2025). Moreover, the DLM architecture allows VidLaDA to employ a full bidirectional attention mechanism without causal masking. Consequently, the aggregation and prediction of any token depends on the entire global context (both unmasked and masked tokens in ${\bm{E}}^{\mathcal{V}}$}, $P$, and $R$), facilitating unconstrained interaction between visual and textual modalities. For detailed mathematical formulations of the forward and reverse diffusion processes, please refer to Appendix B.
@@
Figure 3: Visualization of Attention Patterns. We display the attention score matrices across different decoding steps and layers. The heatmaps reveal two distinct structural properties utilized by our MARS-Cache design: (1) Chunk-wise Locality, visible as diagonal blocks where tokens primarily attend to their temporal neighbors, and (2) Global Anchor Tokens, manifested as prominent vertical bands where specific tokens consistently attract global attention from the entire sequence, regardless of the diffusion step or network depth.
@@ 3.1 Why Bidirectional DLM for Video Understanding?
Subject to non-degenerate encoding assumptions, the AR paradigm inherently limits the cumulative pointwise information capacity, resulting in a strictly lower theoretical upper bound of spatiotemporal understanding compared to a bidirectional decoding paradigm:
$\sum_{t=1}^{T}\sup I({\bm{z}}_{t}^{\text{bidi}};\mathcal{V})>\sum_{t=1}^{T}\sup I({\bm{z}}_{t}^{\text{AR}};\mathcal{V}),$ (1)
where $\mathcal{V}$ denotes the original visual inputs, and ${\bm{z}}^{*}_{t}$ denotes the representation encoded by ViT and LLM at timestep $t$.
@@ 3.1.1 The Impact of Token Spatial Ordering
In ViT, the patch order is conventionally rasterized (top-left to bottom-right). Inspired by (Luo et al., 2025), we use the $\ell_{2}$-norm of the ViT output features as a proxy for the information density of visual tokens. We identify the top-$k$ high-norm tokens and virtually relocate them within the sequence fed to the LLM, but maintaining the original positional IDs in sequence. This design isolates the effect of causal visibility in the decoder from changes in spatial positional encoding. We define a position ratio $r\in[0,1]$, where $r=0$ places high-norm tokens at the start of the visual sequence and $r=1$ pushes them to the end.
As illustrated in Figure 2(a), we evaluate the models on the MME (Yin et al., 2024) benchmark under this shuffling protocol (cf. Appendix G.2). The bidirectional DLM baseline (You et al., 2025) exhibits a near-constant performance profile (red line), demonstrating invariance to the causal position of semantic features.
Conversely, AR baselines (LLaVA-OneVision (Li et al., 2024a), LLaVA-1.5 (Liu et al., 2024b)) display performance degradation, particularly when high-information tokens are shifted towards the middle or end of the sequence. This corroborates Proposition 3.1: AR models over-rely on early tokens as a receptive field to maintain semantic interpretation. When salient features are displaced from the high-visibility start positions to later positions (where visibility is lower), the causal attention mechanism fails to allocate sufficient attention mass to them, treating them as contextually less significant solely due to their temporal index.
@@ 3.1.2 The Impact of Event Position and Sparsity
To verify spatiotemporal robustness in realistic settings beyond artificial single-image shuffling, we evaluate on ReXTime (Chen et al., 2024a) (cf. Appendix G.3), where each question depends on a specific event segment. We bucket samples by the temporal position ratio of the ground-truth event and report accuracy under uniform frame sampling.
@@
Notably, LLaVA-Video (Zhang et al., 2024b) exhibits the most severe instability despite extensive video training compared to LLaVA-OneVision (Li et al., 2024a), suggesting that scaling data alone cannot remove the structural limitations imposed by the early receptive field of AR. This limitation becomes more evident when event evidence is sparse. As shown in Figure 2(c), AR models degrade more sharply as fewer event frames are available, indicating that answer-relevant information of the event is more likely to be dispersed during deep semantic abstraction once it is not continuously reinforced under causal attention. In contrast, the bidirectional DLM baseline maintains a near-flat accuracy profile across event positions (Figure 2(b)) and remains accurate even under sparse event frames (Figure 2(c)). This behavior is consistent with Proposition 3.2: bidirectional decoding allows answer-relevant representations to directly aggregate global spatiotemporal context across all frames. This mechanism enhances understanding efficiency and prevents the irreversible loss of mid-sequence event details during deep semantic abstraction, even under sparse evidence conditions. Furthermore, it inherently eliminates the visibility asymmetry that induces early receptive field in AR decoding (Proposition 3.1).
@@ 3.2 Training Pipeline
Data Composition. To address the scarcity of minute-scale, long-duration video understanding training data, we first curate a dataset centered on 2–30 minute videos. The pipeline involves (1) temporal stratification, (2) instruction synthesis via LLM, (3) text-only bias filtering, and (4) MLLM consistency voting (See Appendix E.1.1 for details).
Multi-Stage Training. We further adopt a multi-stage training strategy for VidLaDA (See Appendix E.1.2 for details), including (1) short-clip temporal pre-training, (2) temporal scaling warm-up, and (3) long-form video expansion.
@@ 3.3 Model Architecture
Bidirectional Diffusion Language Model. The core unit of VidLaDA is the Diffusion Language Model with the full bidirectional attention mechanism. This architecture addresses the limitations of traditional Autoregressive (AR) models to some extent. Standard AR models rely on causal attention, which imposes a strict left-to-right dependency. This structure introduces the asymmetric receptive field, where early visual tokens are structurally prevented from attending to subsequent visual or textual context from user input, disrupting the global topology of the vision (See Section 3.1). In contrast, VidLaDA leverages a full bidirectional attention mechanism. This design eliminates the causal constraint, allowing every token, whether visual or textual, to attend to the holistic sequence simultaneously. This ensures that the vision encoder and language model are deeply coupled, preserving the integrity of dynamic video information.
@@ 4.1 Empirical Observations
We analyze the internal state evolution during the decoding process and identify four distinct observations: (1) Chunk-wise Locality with Global Anchor Tokens. As shown in Figure 3, inter-frame interactions primarily exhibit local dependencies. However, a specific subset of anchor tokens consistently attracts high attention scores globally. Crucially, we observe that the spatial positions of these anchor tokens remain stable across decoding steps, indicating that the global information hubs are determined early in the inference process. Furthermore, we observe a hierarchical inclusion property: the set of anchor tokens in deeper layers is a subset of those in shallow layers ($\mathcal{I}_{\text{deep}}\subset\mathcal{I}_{\text{shallow}}$), adhering to a dense-to-sparse pyramid structure. (2) Modality-Dependent Drift. As shown in Figure 5, there is a notable disparity in hidden state semantic evolution between modalities. For the part of user inputs, text hidden states exhibit higher temporal drift across steps compared to vision tokens. (3) Depth-Dependent Stability Variance. As shown in Figure 5, hidden state drift is not uniform across network depth. Shallow layers exhibit high stability, while deep layers show significant volatility. (4) Progressive Attention Sparsity. As shown in Figure 6, the attention distribution evolves through the network depth, transitioning from a near-uniform distribution in shallow layers to a highly peaked, sparse distribution in deep layers.
@@
Table 1: Comparison with SOTA models across comprehensive benchmarks. The models are categorized into AR and DLM baselines. #S: Model Size (Parameters). #F: Number of input frames. Models marked with * denote the reproduced results.